OpenStudy (alyssajobug):

Help with Algebra 2? [−1 −2]^3 [ 1 0] Can I do that?

OpenStudy (kc_kennylau):

\[\left(\begin{matrix}-1&-2\\1&0\end{matrix}\right)\left(\begin{matrix}-1&-2\\1&0\end{matrix}\right)\left(\begin{matrix}-1&-2\\1&0\end{matrix}\right)\]\[=\left(\begin{matrix}-1\times(-1)+(-2)\times1&-1\times(-2)+(-2)\times0\\1\times(-1)+0\times1&1\times(-2)+0\times0\end{matrix}\right)\left(\begin{matrix}-1&-2\\1&0\end{matrix}\right)\]\[=\left(\begin{matrix}-1&2\\-1&-2\end{matrix}\right)\left(\begin{matrix}-1&-2\\1&0\end{matrix}\right)\]\[=\left(\begin{matrix}-1\times(-1)+2\times1&-1\times(-2)+2\times0\\-1\times(-1)+(-2)\times1&-1\times(-2)+(-2)\times0\end{matrix}\right)\]\[=\left(\begin{matrix}3&4\\-1&2\end{matrix}\right)\] Remember the golden rule: \[\left(\begin{matrix}a&b\\c&d\end{matrix}\right)\left(\begin{matrix}e&f\\g&h\end{matrix}\right)=\left(\begin{matrix}ae+bg&af+bh\\ce+dg&cf+dh\end{matrix}\right)\]

OpenStudy (alyssajobug):

these are my options

OpenStudy (alyssajobug):

Help?!?!?!?!?!?!?!

OpenStudy (anonymous):

look what @kc_kennylau did

OpenStudy (alyssajobug):

It isnt an option tho...

OpenStudy (alyssajobug):

@helpme1.2

OpenStudy (anonymous):

KC-math is a little off at line 4

OpenStudy (anonymous):